A liter of D5/0.45 NS with 30 mEq KCl is to infuse over 12 hours. How many mL / hr will the nurse set on the IV pump? ________ mL / hr
Answer #1

Given

Fluid:D5/0.45NS=1litre (1000ml)

Medication :30mEq KCL

Time:12 hour

Flow rate =Total volume ×15÷Total time of transfusion ×60

=1000×15÷12×60

=20.8 drops /min or 21 drops /min

When 12 hours =1000ml

Then 1 hour =??

Cross multiply and divide

=1000×1÷12

=83.3 or 83ml/hr

The nurse has to set at 83ml/hr